OpenStudy (anonymous):

20% of what number is 47

OpenStudy (arbershabani97):

20/100*x=47 x=47/(2/10) x=9.4

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

47/.2 = 235

OpenStudy (anonymous):

estimate a 15% tip for a 26.80 meal

OpenStudy (anonymous):

ok

jigglypuff314 (jigglypuff314):

15% -> 0.15 so 0.15 * 26.8 = (tip amount)

OpenStudy (anonymous):

thx

jigglypuff314 (jigglypuff314):

yw